# Inverse analysis of granular flows using differentiable graph neural network simulator

### v2

@@ Abstract
Inverse problems in granular flows, such as landslides and debris flows, involve estimating material parameters or boundary conditions based on target runout profile. Traditional high-fidelity simulators for these inverse problems are computationally demanding, restricting the number of simulations possible. Additionally, their non-differentiable nature makes gradient-based optimization methods, known for their efficiency in high-dimensional problems, inapplicable. While machine learning-based surrogate models offer computational efficiency and differentiability, they often struggle to generalize beyond their training data due to their reliance on low-dimensional input-output mappings that fail to capture the complete physics of granular flows. We propose a novel differentiable graph neural network simulator (GNS) by combining reverse mode automatic differentiation of graph neural networks with gradient-based optimization for solving inverse problems. GNS learns the dynamics of granular flow by representing the system as a graph and predicts the evolution of the graph at the next time step, given the current state. The differentiable GNS shows optimization capabilities beyond the training data. We demonstrate the effectiveness of our method for inverse estimation across single and multi-parameter optimization problems, including evaluating material properties and boundary conditions for a target runout distance and designing baffle locations to limit a landslide runout. Our proposed differentiable GNS framework offers an orders of magnitude faster solution to these inverse problems than the conventional finite difference approach to gradient-based optimization.
@@ 1 Introduction
In geotechnical engineering, addressing optimization problems is crucial, such as the inverse modeling of landslides to infer material properties and design protective structures. Inverse or back analysis is crucial in analyzing hazards such as landslides, which involves accurately deducing material properties from landslide runout patterns. Traditionally, this process has been approached by iteratively adjusting material properties to closely match observed results, often resulting in oversimplified, integer-valued material parameters. Design problems focus on optimizing geometries, such as the location and geometry of protective structures, to effectively mitigate and control granular flows (Liu et al., 2023; Babu and Basha, 2008). Addressing inverse and design optimization challenges is critical to developing robust models for managing landslide risks (Calvello et al., 2017; Cuomo et al., 2015; Abraham et al., 2021).
@@
Sampling-based methods like grid search (Ensor and Glynn, 1997), cross-entropy (Ho and Yang, 2010), and Bayesian optimization (Frazier, 2018) often require a significant number of forward simulations, especially as the number of parameters increases. In contrast, gradient-based optimization utilizes derivatives of the objective function to navigate the parameter space efficiently. This approach often leads to faster convergence with fewer simulations but requires a differentiable objective function. However, traditional simulators are not differentiable. While finite difference methods can approximate the derivatives, they require a smaller increment size and are susceptible to numerical errors.
@@
Recent advancements in learned physics simulators offer a promising solution to these limitations (Battaglia et al., 2016, 2018; Sanchez-Gonzalez et al., 2020). Graph neural network-based simulators (GNNs) are one such learned physics simulator that represents the domain as a graph with vertices and edges and learns the local interactions between the vertices, which are critical in governing the dynamics of physical systems. By learning the interaction law, GNS demonstrates high predictive accuracy and generalization ability in modeling granular flows beyond the training data (Choi and Kumar, 2023a). Since GNS is based on neural network foundations, it is inherently differentiable through automatic differentiation (AD), making it suitable for gradient-based optimization (Allen et al., 2022; Zhao et al., 2022).
@@
A key challenge in applying AD to GNS, especially in simulations with extended temporal scales, is the extensive GPU memory requirement (Zhao et al., 2022). AD necessitates tracking all intermediate computations of GNS over the entire simulation for gradient evaluation. To mitigate this issue, we employ gradient checkpointing proposed by Chen et al. (2016). This approach significantly reduces the memory demand by strategically storing only key intermediate computational steps during gradient evaluation. While this technique conserves memory, it requires re-evaluating certain forward computations during the backward pass for gradient evaluation. This hybrid approach allows monitoring the gradients across thousands of particle trajectories over prolonged simulation durations without overwhelming the GPU memory, albeit with an increased computational burden.
We introduce a novel framework for inverse analysis in granular flows, leveraging the automatic differentiation capabilities of GNS with gradient-based optimization. We demonstrate our gradient-based optimization framework on inverse problems and design applications in granular flows. In the following sections, we describe the details of the proposed differentiable AD-GNS and gradient-based optimization approaches for solving inverse problems.
@@ 3.1.1 Training data
We train the GNS model for the two-dimensional domain (fig. 1a and fig. 1b) on the Flow2D dataset. It includes 385 square-shaped granular mass flow trajectories in a two-dimensional box boundary. We use five different friction angles ($\phi=15\degree,22.5\degree,30\degree,37.5\degree,45\degree$), where each friction angle has 77 simulation datasets. Each simulation has different initial configurations regarding the size of the square granular mass, position, and velocity. We use the explicit time integration method using CB-Geo MPM code (Kumar et al., 2019) to generate the training data.
We train the 3D GNS (see fig. 1c) on the Obstacle3D dataset. It includes 1000 trajectories of granular mass interacting with barriers in a three-dimensional box boundary. The initial geometry of the cuboid-shaped granular mass varies from 0.25 to 0.80 $m$ and is excited with different initial velocities. For each simulation, the granular mass interacts with one to three cuboid-shaped barriers whose width and length vary from 0.10 to 0.13 $m$ with a height of 0.3 $m$ in $1.0\times 1.0\times 1.0\ m$ domain. We use the explicit time integration method using Moving Least Squares (MLS)-MPM code (Hu et al., 2018) to generate the training data. See A for the forward simulation performance of GNS.
@@ 3.2 Differentiable GNS
The next step is backward pass (fig. 4b), known as reverse mode or backpropagation, which is the core process of reverse-mode AD to compute the gradient with respect to the inputs of the function ($x_{1}$ and $x_{2}$). This phase starts from the output of the computational graph ($y$) and works backward. As the process moves backward through the graph, reverse-mode AD computes what is known as a local derivative—the derivative of that node’s operation with respect to its immediate input. In fig. 4b, this corresponds to $\partial{y}/\partial{x_{1}}$, $\partial{y}/\partial{w_{1}}$, $\partial{w_{1}}/\partial{x_{1}}$, and $\partial{w_{1}}/\partial{x_{2}}$, denoted in the dashed line boxes. By representing the simulation as a computational graph, we can pre-compute the derivatives, thus improving computational efficiency in gradient evaluation. We employ the chain rule, which combines every local derivative at each node from the output to the input node, returning the gradient of output with respect to the graph’s input ($\partial{y}/\partial{x_{1}}$ and $\partial{y}/\partial{x_{2}}$). In fig. 4b, this process is described below the nodes $x_{1}$ and $x_{2}$.
@@ 3.3 Gradient-based optimization
We use adaptive movement estimation algorithm (ADAM) (Kingma and Ba, 2014) for high-dimensional parameter optimization. ADAM incorporates a dynamic adaptive learning rate for each parameter using the moments of the gradients. ADAM accelerates the parameter updates in the right direction, thus avoiding local minima and adapting the learning rate for each parameter, making it less sensitive to the scale of the gradients. This adaptive learning rate is particularly effective optimization in high-dimensional parameter spaces, which entails sparse and varied gradient scales. This adaptive learning makes ADAM converge quicker and improves performance in solving multi-parameter inverse problems. We refer readers to Kingma and Ba (2014) for more technical details.
@@ 3.3.1 Gradient checkpointing
Reverse-mode AD requires significant memory for large-scale neural networks (Zhao et al., 2022). AD necessitates storing intermediate variables in the computational graph during the forward pass. These intermediate variables should be retained during the backward pass to apply the chain rule in computing gradients. The computational graph grows in size for large-scale neural networks by adding more layers and parameters. Therefore, the backpropagation requires substantial memory to retain all the intermediate variables in the increasingly large computational graph.
@@
We integrate gradient checkpointing in our GNS-based optimization. We include a checkpoint at each timestep during rollout, which allows the implementation of reverse-mode AD across extended simulation timesteps within existing memory constraints without the need for proportional increases in hardware capability. Consequently, we achieve gradient computation over hundreds of timesteps, significantly improving from the vanilla AD-GNS of 3 or 4 timesteps for the same GPU capacity.

### v3

@@ Abstract
Inverse problems in granular flows, such as landslides and debris flows, involve estimating material parameters or boundary conditions based on a target runout profile. Traditional high-fidelity simulators for these inverse problems are computationally expensive, restricting the number of possible simulations. These simulators are also non-differentiable, making efficient gradient-based optimization methods for high-dimensional spaces inapplicable. Machine learning-based surrogate models offer computational efficiency and differentiability. However, they often struggle to generalize beyond their training data due to relying on low-dimensional input-output mappings that fail to capture the complete physics of granular flows. We propose a novel differentiable graph neural network simulator (GNS) that combines reverse mode automatic differentiation of graph neural networks with gradient-based optimization for solving inverse problems. GNS learns the dynamics of granular flow by representing the system as a graph and predicts the evolution of the graph at the next timestep, given the current state. The differentiable GNS demonstrates optimization capabilities beyond the training data. We demonstrate the effectiveness of our method for inverse estimation across single and multi-parameter optimization problems, including evaluating material properties and boundary conditions for a target runout distance and designing baffle locations to limit a landslide runout. Our proposed differentiable GNS framework solves inverse problems with orders of magnitude faster convergence than the conventional gradient-based optimization approach using finite difference.
@@
journal: Computers and Geotechnics
[…]
[color=purple, name=Yongjin Choi]YC
@@ 1 Introduction
In geotechnical engineering, addressing optimization problems, such as the inverse modeling of landslides to infer material properties from target runout distances and design protective structures, is essential. Traditionally, this inverse optimization process has been approached by iteratively adjusting material properties to closely match observed results, often resulting in oversimplified, integer-valued material parameters. Design problems focus on optimizing geometries, such as the location and geometry of protective structures, to effectively mitigate and control granular flows (Liu et al., 2023; Babu and Basha, 2008). Addressing inverse and design optimization challenges is critical to developing robust models for managing landslide risks (Calvello et al., 2017; Cuomo et al., 2015; Abraham et al., 2021).
@@
Sampling-based methods like grid search (Ensor and Glynn, 1997), cross-entropy (Ho and Yang, 2010), and Bayesian optimization (Frazier, 2018) often require a significant number of forward simulations, especially as the number of parameters increases. In contrast, gradient-based optimization utilizes derivatives of the objective function to navigate the parameter space efficiently. This approach often leads to faster convergence with fewer simulations but requires the gradient of the objective function. However, traditional simulators are not differentiable. While finite difference methods can approximate the derivatives, they require a smaller increment size and are susceptible to numerical errors.
@@
Recent advancements in learned physics simulators offer a promising solution to these limitations (Battaglia et al., 2016, 2018; Sanchez-Gonzalez et al., 2020). Graph neural network (GNN)-based simulators (GNSs) are one such learned physics simulator that represents the domain as a graph with vertices and edges and learns the local interactions between the vertices, which are critical in governing the dynamics of physical systems. By learning the interaction law, GNS demonstrates high predictive accuracy and generalization ability in modeling granular flows beyond the training data (Choi and Kumar, 2023a). Since GNS is based on neural network foundations, it is inherently differentiable through automatic differentiation (AD), making it suitable for gradient-based optimization (Allen et al., 2022; Zhao et al., 2022).
@@
A key challenge in applying AD to GNS, especially in simulations with extended temporal scales, is the extensive GPU memory requirement (Zhao et al., 2022). AD necessitates tracking all intermediate computations of GNS over the entire simulation for gradient evaluation. To mitigate this issue, we employ gradient checkpointing proposed by Chen et al. (2016). This approach significantly reduces the memory demand by strategically storing only key intermediate computational steps during gradient evaluation. While this technique conserves memory, it requires re-evaluating certain forward computations during the backward pass for gradient evaluation. This hybrid approach allows for monitoring gradients across thousands of particle trajectories over prolonged simulation durations without overwhelming GPU memory, albeit with an increased computational burden.
We introduce a novel framework for inverse analysis in granular flows, leveraging the automatic differentiation capabilities of GNS with gradient-based optimization. We demonstrate our gradient-based optimization framework on inverse problems and design applications in granular flows. The overview of the research is shown in fig. 1. We first create granular flow trajectories using the material point method (MPM) (Soga et al., 2016; Kumar et al., 2019; Hu et al., 2018). The trajectories are stored as NPZ files (compressed NumPy file format) and are used to train GNS. We then use GNS as a surrogate to accelerate the forward evaluation of granular flow runouts. To solve the inverse problem, we leverage the reverse-mode Automatic Differentiation of GNS (AD-GNS) nature of GNS to calculate the derivatives of the target with respect to the input parameters. Finally, we use the gradients to optimize the parameters of interest based on a target iteratively. In the following sections, we describe the details of the proposed differentiable AD-GNS and gradient-based optimization approaches for solving inverse problems.
Figure 1: Overview of the research.
@@ 3.1 Graph neural network-based forward simulator (GNS)
We leverage PyTorch’s Distributed Data Parallelism (DDP) framework to enable multiple GPU parallelization. This approach demonstrates almost linear scaling performance (Kumar and Vantassel, 2022; Choi and Kumar, 2023b). By employing DDP, we can distribute the workload associated with large graph structures across multiple GPUs with batching, thus facilitating the operation of GNS with a distributed computational burden without compromising the speed.
@@ 3.1.1 Training data
We train the GNS model for the two-dimensional domain (fig. 2a and fig. 2b) on the Flow2D dataset. It includes 385 square-shaped granular mass flow trajectories in a two-dimensional box boundary. The appendix section provides a visual example in fig. 1. We use five different friction angles ($\phi=15\degree,22.5\degree,30\degree,37.5\degree,45\degree$), where each friction angle has 77 simulation datasets. Each simulation has different initial configurations regarding the size of the square granular mass, position, and velocity. We use the explicit time integration method using CB-Geo MPM code (Kumar et al., 2019) to generate the training data.
We train the 3D GNS (see fig. 2c) on the Obstacle3D dataset. It includes 1000 trajectories of granular mass interacting with barriers in a three-dimensional box boundary. The appendix section provides a visual example in fig. 2. The initial geometry of the cuboid-shaped granular mass varies from 0.25 to 0.80 $m$ and is excited with different initial velocities. For each simulation, the granular mass interacts with one to three cuboid-shaped barriers whose width and length vary from 0.10 to 0.13 $m$ with a height of 0.3 $m$ in the $1.0\times 1.0\times 1.0\ m$ domain. We use the explicit time integration method using Moving Least Squares (MLS)-MPM code (Taichi MPM) (Hu et al., 2018) to generate the training data. See A for the forward simulation performance of GNS.
We use two different MPM software, CB-Geo MPM (2D simulations) and Taichi MPM (3D simulations), to get the best of both worlds. While CB-Geo MPM can reasonably simulate 2D granular flow, a 3D MPM code with rigid body dynamics was needed to model the barrier problems – hence, we use Taichi MPM code.
@@ 3.2 Differentiable GNS
The next step is backward pass (fig. 5b), known as reverse mode or backpropagation, which is the core process of reverse-mode AD to compute the gradient with respect to the inputs of the function ($x_{1}$ and $x_{2}$). This phase starts from the output of the computational graph ($y$) and works backward. As the process moves backward through the graph, reverse-mode AD computes what is known as a local derivative—the derivative of that node’s operation with respect to its immediate input. In fig. 5b, this corresponds to $\partial{y}/\partial{w_{4}}$, $\partial{w_{4}}/\partial{w_{3}}$, $\partial{w_{3}}/\partial{w_{1}}$, and $\partial{w_{4}}/\partial{w_{2}}$, denoted beside the gray dashed edges. By representing the simulation as a computational graph, we can pre-compute the derivatives, thus improving computational efficiency in gradient evaluation. We employ the chain rule, which combines every local derivative at each node from the output to the input node, returning the gradient of output with respect to the graph’s input ($\partial{y}/\partial{x_{1}}$, $\partial{y}/\partial{x_{2}}$). In fig. 5b, this process is described below the nodes $x_{1}$ and $x_{2}$.
@@
The forward pass of the GNS is described as:
$\boldsymbol{X}_{t+1}=GNS(\boldsymbol{X}_{t})\quad\mathrm{for}\ t\in\{0,1,\ldots,k-1\}$ (1)
where $\boldsymbol{X}_{t}$ is the current state of the system at time $t$, and GNS is the learned function approximator that predicts the next state $\boldsymbol{X}_{t+1}$ given the current state $\boldsymbol{X}_{t}$. The forward pass involves predicting the trajectory for $k$ timesteps to simulate the evolution of the system from the initial state $\boldsymbol{X}_{0}$ to $\boldsymbol{X}_{k}$.
For the inverse analysis, we define an objective function (loss) $J_{\boldsymbol{\Theta}}$ that needs to be minimized with respect to the parameters ${\boldsymbol{\Theta}}$. The objective function depends on the output of GNS after rolling out for $k$ timesteps, starting from the initial state $\boldsymbol{X}_{0}$.
$J_{\boldsymbol{\Theta}}=J(\boldsymbol{X}_{k})$ (2)
where $\boldsymbol{X}_{k}$ is obtained by recursively applying $GNS()$ for $k$ timesteps:
$\boldsymbol{X}_{1}$ $=GNS(\boldsymbol{X}_{0};{\boldsymbol{\Theta}})$
$\boldsymbol{X}_{2}$ $=GNS(\boldsymbol{X}_{1};{\boldsymbol{\Theta}})$
$\vdots$
$\boldsymbol{X}_{k}$ $=GNS(\boldsymbol{X}_{k-1};{\boldsymbol{\Theta}})$
We can compactly represent this recursive forward pass as:
$\boldsymbol{X}_{k}=GNS^{k}(\boldsymbol{X}_{0};{\boldsymbol{\Theta}})$ (3)
where $GNS^{k}$ denotes applying $GNS()$ recursively for $k$ timesteps.
To compute the gradient of $J_{\boldsymbol{\Theta}}$ with respect to ${\boldsymbol{\Theta}}$ using reverse mode AD, we construct the computational graph for the entire forward pass $GNS^{k}$ and then apply the chain rule:
$\nabla_{\boldsymbol{\Theta}}J_{\boldsymbol{\Theta}}=\frac{\partial J}{\partial\boldsymbol{X}_{k}}\frac{\partial\boldsymbol{X}_{k}}{\partial\boldsymbol{\Theta}}$ (4)
where $\frac{\partial\boldsymbol{X}_{k}}{\partial\boldsymbol{\Theta}}$ is computed by recursively applying the chain rule through the entire trajectory:
$\frac{\partial\boldsymbol{X}_{k}}{\partial\boldsymbol{\Theta}}=\frac{\partial\boldsymbol{X}_{k}}{\partial\boldsymbol{X}_{k-1}}\ldots\frac{\partial\boldsymbol{X}_{1}}{\partial\boldsymbol{X}_{0}}\frac{\partial\boldsymbol{X}_{0}}{\partial\boldsymbol{\Theta}}$ (5)
This reverse mode AD computes the exact gradients $\nabla_{\boldsymbol{\Theta}}J_{\boldsymbol{\Theta}}$ by propagating the gradients from the final state $\boldsymbol{X}_{k}$ back to the initial state $\boldsymbol{X}_{0}$ and then to the parameters $\boldsymbol{\Theta}$, while leveraging the intermediate states $\boldsymbol{X}_{1},\boldsymbol{X}_{2},\cdots,\boldsymbol{X}_{k-1}$ computed during the forward pass.
Note that storing all these intermediate states for long trajectories (large $k$) can be memory-intensive, which is where gradient checkpointing (section 3.3.1) is employed to reduce the memory footprint by selectively storing and recomputing intermediate states.
@@ 3.3 Gradient-based optimization
The learning rate or step size in an optimization algorithm determines the increment at each iteration while traversing toward a minimum of a loss function; this learning rate or step size $\eta$, multiplied by the gradient $\nabla_{\boldsymbol{\Theta}}J_{\boldsymbol{\Theta}}$, controls the aggressiveness of parameter updates in a gradient descent optimization of multi-variable function $J_{\boldsymbol{\Theta}}$ (see eq. 6). A high learning rate can lead to instability in the optimization process, with large steps causing overshooting of the minimum or even divergence. Conversely, a small learning rate will result in slow convergence, as the model takes minuscule steps and requires excessive iterations to reach a solution. Finding the optimal learning rate is largely experimental. We start with a small learning rate (e.g., 0.001), observe the parameter updates in the initial iterations, and manually vary the learning rate to achieve convergence in a reasonable time. We can also employ adaptive learning rates or learning rate schedulers that decrease the rate over time for stability. Alternatively, adaptive optimizers like adaptive movement estimation algorithm (ADAM) (Kingma and Ba, 2014) or root mean squared propagation (RMSprop) automate some of the learning rate adjustment process.
@@
We use the adaptive movement estimation algorithm (ADAM) (Kingma and Ba, 2014) for high-dimensional parameter optimization. ADAM incorporates a dynamic adaptive learning rate for each parameter using the moments of the gradients. ADAM accelerates the parameter updates in the right direction, thus avoiding local minima and adapting the learning rate for each parameter, making it less sensitive to the scale of the gradients.
Specifically, the parameters are updated as:
$\boldsymbol{\Theta}_{i}=\boldsymbol{\Theta}_{i-1}-\frac{\eta}{\sqrt{\hat{v}_{i}}+\epsilon}\hat{m}_{i}$ (7)
where $\boldsymbol{\Theta}$ is the parameter set with the optimization step $i$, and $\epsilon$ is a small scalar for numerical stability. The first and second moments ($m_{i}$ and $v_{i}$) of the gradients are updated as:
$m_{i}$ $=\beta_{1}m_{i-1}+(1-\beta_{1})g_{i}$ (8)
$v_{i}$ $=\beta_{2}v_{i-1}+(1-\beta_{2})g_{i}^{2}$ (9)
where $g_{i}$ represents the gradient at step $i$, $\beta_{1}$ and $\beta_{2}$ are decay rates that control the exponential decay of these moving averages. Early in the training process, the variables $m_{i}$ and $v_{i}$ are initially biased toward zero because they start from zero. To address this bias, a correction is applied as:
$\hat{m}_{i}$ $=\frac{m_{i}}{1-\beta_{1}^{i}}$ (10)
$\hat{v}_{i}$ $=\frac{v_{i}}{1-\beta_{2}^{i}}$ (11)
This adaptive learning rate is particularly effective optimization in high-dimensional parameter spaces, which entails sparse and varied gradient scales. ADAM adapts the learning rate for each parameter based on its historical gradients, which helps in scenarios where different parameters have different scales or degrees of sparsity. This adaptive learning makes ADAM converge quicker and improves performance in solving multi-parameter inverse problems. We refer readers to Kingma and Ba (2014) for more technical details.
@@ 3.3.1 Gradient checkpointing
Reverse-mode AD requires significant memory for large-scale neural networks (Zhao et al., 2022). AD necessitates storing intermediate variables in the computational graph during the forward pass. These intermediate variables should be retained during the backward pass to apply the chain rule in computing gradients. The computational graph grows for large-scale neural networks by adding more layers and parameters. Therefore, the backpropagation requires substantial memory to retain all the intermediate variables in the increasingly large computational graph.
@@
Let’s consider the forward pass of GNS over $k$ timesteps:
$\boldsymbol{X}_{0}\rightarrow\boldsymbol{X}_{1}\rightarrow\boldsymbol{X}_{2}\cdots\rightarrow\boldsymbol{X}_{k}$ (12)
During the forward pass, instead of storing all the intermediate states $\boldsymbol{X}_{0},\boldsymbol{X}_{1},\boldsymbol{X}_{2},\cdots,\boldsymbol{X}_{k-1}$, gradient checkpointing selectively stores only a subset of these states, called checkpoints, at specific timesteps $C\in\{0,c_{1},c_{2},\cdots,c_{m},k\}$ as:
$\boldsymbol{X}_{i}=\text{GNS}(\boldsymbol{X}_{i-1});\quad\left\{\begin{array}[]{ll}\text{store}&\leftarrow\boldsymbol{X}_{i}\quad\forall i\in C,\\ \text{discard}&\leftarrow\boldsymbol{X}_{i}\quad\forall i\notin C\end{array}\right.$ (13)
During the backward pass, those intermediate states between $\boldsymbol{X}_{i-1}$ and $\boldsymbol{X}_{i}$ $\forall i\notin C$ are recomputed to evaluate the local derivatives. The reverse mode AD with gradient checkpoints is computed as follows:
  1. 1.
    Initial gradient computation at final state: Start by computing the gradient of the loss function $J$ with respect to the final state $\boldsymbol{X}_{k}$: $\frac{\partial J}{\partial\boldsymbol{X}_{k}}$ This gradient is derived from the final output of the GNS model, based on how the loss $J$ is affected by the last computed state $\boldsymbol{X}_{k}$.
  2. 2.
    Backpropagate through GNS from $\boldsymbol{X}_{k}$ to last checkpoint $\boldsymbol{X}_{c_{m}}$:
    Use the gradient from the final state to backpropagate through the operations performed from the last checkpoint to the final state:
    $\frac{\partial J}{\partial\boldsymbol{X}_{c_{m}}}=\left(\frac{\partial J}{\partial\boldsymbol{X}_{k}}\right)\cdot\left(\frac{\partial\boldsymbol{X}_{k}}{\partial\boldsymbol{X}_{c_{m}}}\right)$
    Here, $\frac{\partial\boldsymbol{X}_{k}}{\partial\boldsymbol{X}_{c_{m}}}$ represents the gradient of the output at step $k$ with respect to the state at the last checkpoint $c_{m}$, encompassing all intermediate transformations.
  3. 3.
    Backpropagate successively through each checkpoint:
    For each checkpoint $c_{m}$, starting from the last stored checkpoint $c_{m}$ and moving backward through earlier checkpoints $c_{m-1},c_{m-2},\ldots$, compute:
    $\frac{\partial J}{\partial\boldsymbol{X}_{c_{(m-1)}}}=\left(\frac{\partial J}{\partial\boldsymbol{X}_{c_{m}}}\right)\cdot\left(\frac{\partial\boldsymbol{X}_{c_{m}}}{\partial\boldsymbol{X}_{c_{(m-1)}}}\right)$
    Each term $\frac{\partial\boldsymbol{X}_{c_{m}}}{\partial\boldsymbol{X}_{c_{(m-1)}}}$ captures the gradient of the checkpointed state with respect to its preceding checkpoint, reflecting the operations of GNS executed between these points.
  4. 4.
    Recompute non-stored states as needed:
    Assuming a non-checkpointed intermediate state $i$ between checkpoints $c_{(m-1)}$ and $c_{(m-2)}$, where $c_{(m-2)}<i<c_{(m-1)}$ and $\boldsymbol{X}_{i}$ is not stored during the forward pass. We handle the backward pass as:
    1. (a)
      Recompute intermediate states: If $\boldsymbol{X}_{i}$ is an intermediate state that was not stored, recomputation from the last stored state $\boldsymbol{X}_{c_{(m-2)}}$ is necessary:
      $\boldsymbol{X}_{i}=\text{GNS}^{i\leftarrow c_{(m-2)}}(\boldsymbol{X}_{c_{(m-2)}})$
      This step involves executing the forward function GNS iteratively from $c_{(m-2)}$ to $i$, reconstructing each $\boldsymbol{X}_{i}$ sequentially up to $\boldsymbol{X}_{c_{(m-1)}}$.
    2. (b)
      Backpropagate gradient through $\boldsymbol{X}_{i}$: With $\boldsymbol{X}_{i}$ recomputed, the gradient from $\boldsymbol{X}_{c_{(m-1)}}$ is propagated back:
      $\frac{\partial J}{\partial\boldsymbol{X}_{i}}=\left(\frac{\partial J}{\partial\boldsymbol{X}_{c_{(m-1)}}}\right)\cdot\left(\frac{\partial\boldsymbol{X}_{c_{(m-1)}}}{\partial\boldsymbol{X}_{i}}\right)$
      We compute the gradients for each $\boldsymbol{X}_{j}$ (for $j=k$ to $c_{(m-1)}$) during this backpropagation process.
    3. (c)
      Gradient at checkpoint $c_{(m-2)}$: We continue the gradient propagation back to $\boldsymbol{X}_{c_{(m-2)}}$:
      $\frac{\partial J}{\partial\boldsymbol{X}_{c_{(m-2)}}}=\left(\frac{\partial J}{\partial\boldsymbol{X}_{i}}\right)\cdot\left(\frac{\partial\boldsymbol{X}_{i}}{\partial\boldsymbol{X}_{c_{(m-2)}}}\right)$
      This step involves recalculating each partial derivative $\partial\boldsymbol{X}_{j}/\partial\boldsymbol{X}_{j-1}$ during the recomputation phase.
  5. 5.
    Finalize gradients for initial state and parameters:
    Once backpropagation reaches the first checkpoint $\boldsymbol{X}_{c_{1}}$, complete the gradient computation for the initial state and model parameters:
    $\frac{\partial J}{\partial\boldsymbol{X}_{0}}=\left(\frac{\partial J}{\partial\boldsymbol{X}_{c_{1}}}\right)\cdot\left(\frac{\partial\boldsymbol{X}_{c_{1}}}{\partial\boldsymbol{X}_{0}}\right)$
    $\frac{\partial J}{\partial\boldsymbol{\Theta}}=\left(\frac{\partial J}{\partial\boldsymbol{X}_{c_{1}}}\right)\cdot\left(\frac{\partial\boldsymbol{X}_{c_{1}}}{\partial\boldsymbol{\Theta}}\right)$
    These calculations utilize the gradients at the first checkpoint to determine how initial conditions and parameters should be adjusted to minimize the loss $J$.
By selectively storing and recomputing intermediate states, gradient checkpointing reduces the peak memory requirement during backpropagation at the cost of some additional computations during the backward pass. Choosing checkpoint locations $C$ is crucial for balancing memory savings and computational overhead. A common strategy is to place checkpoints at regular intervals (e.g., every few timesteps) or at specific points in the simulation where the computational graph is expected to be particularly large or complex. Using gradient checkpointing, we can differentiate through the entire simulation.